(4)/(x^2-4)+(2x)/(x-2)+2/(x+2)=0
Answer by stanbon(75887)

(4)/(x^2-4)+(2x)/(x-2)+2/(x+2)=0
------------
Multiply thru by x^2-4 to get:
4 + 2x(x+2) + 2(x-2) = 0
---
4 + 2x^2+4x + 2x-4 = 0
2x^2 + 6x -2 = 0
-----
x^2 + 3x -1 = 0
---
x = [-3 +- sqrt(9 -4*-1)]/2
---
x = [-3 +- sqrt(13)]/2
=============================
Cheers,
stan H.